MODS1 AGw Commissioning at LDG: 17 November 2010 UTC

Observer(s): R. Pogge and B. Shappee (OSU), M. Pedani and M. Wagner (LBTO)
Telescope Operator: G. Bechetti (LBTO)
Active Optics Support: J. Hill (LBTO) and D. Miller (LBTO remote from Tucson)
Telescope Support: J. Little (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

Clear and beautiful, but windy at sunset. At 0 UTC the weather is T=0C, 59% humidity trending upward the previous 4 hours, wind 12-18m/s out of the WNW with some strong gusts. High winds kept us from opening until 6 UT.

Seeing was 1-arcsec at best, but highly variable, and winds stayed above 10m/s for most of the night, but never again caused us to close.

Worked through transformation acquisition during a good seeing interval, but had numerous problems.

Characterized a feature of the AGw stage that was causing some preset failures due to the stage Y-axis running out of travel on the move-star-to-WFS stage offset at the extreme limits of Y travel. Note that some presets were outside the nominal range for WFS, but valid for initial Guide Probe placement. Drawing in the lower boundary of the patrol field by a few arcsec should mitigate the problem.

Lost a little time (30m increments) due to a glitch in the elevation drive and the need to reboot the AGw guide camera AzCam computer.

Preparations

LN2 filled at 1630, still liquid in the reservoirs.

Details

Startup

0100 - winds too strong to open anything but the vent door to try to get some of the warm air out. Gusts are up to 20m/sec

0250 - Now getting frequent gusts above 22m/s, vent doors being closed for the time being.

IMCS Telescope Scan

0300 - With the wind above limits consisently, we're using the closed dome time to finish the IMCS telescope scan from Nov 14.

This scan set is the 45-degree compass points in position angle (raw rotator angles 255, 345, 165, and 75) in elevations 90 to 30 in 15-degrees steps.

0440 - wind starting to trend down... One more scan sequence to finish.

0500 - and then back up to the limit. No breaks so far.

IMCS telescope scan completed.

Opening

0540 - Wind has been below the limit long enough to want to try opening up

0545 - Elevation drive stalled at about El=15-deg while driving towards horizon preparatory to opening the doors. Technical personnel having a look...

0613 - We're back, nobody found anything obviously wrong. The proposal is to blame the moths.

0616 - Pointing target to start and check collimation: BS910 (02h33/06d25)

problems finding the pointing star, much fussing about...

0705 - Found the star, some kind of glitch on the GCS and stage? I took a science image with the red channel, star was near dead center, then reset the stage and drove it by hand to field center.

New pointing constants:
   IE=15
   CA=19

0708 - Setting on the star for collimation adjustment

Seeing is about 1 arcsec, but the mirrors are still warmer than the air. Wind is now 13m/sec, gusts up to 15m/sec, humidity 60% and flopping around, Temperature is -2C and falling (but flatter than earlier). Might be OK once the telescope equilibrates.

AGw Coordinate Transformation

Stone C field prepared by Doug Miller. Seeing 1.1-arcsec, centered on "star 0".

0725 - Using 10s exposure with SDSS r filter on the red channel (blue channel is idle for this task), but...

The IDL script failed with an error (tried to send something to the IIF it didn't like, a typo in the script?). The seeing is puffing up ridiculously, so we're not in a great hurry. POSANG was changed to PA in the calls to iif_preset.pro.

0732 - Trying again...works, but the seeing is dreadful...

WFS images 7, 8, and 9 are not well centered within the analysis grid, off by 2 columns of spots.
This was caused by the actual xy probe positions being reported as 0,0 in the header rather than the actual values.

On star 18, collimation is not converging, we're stuck here...

0740 - Preset for star 17, the preset nearly failed. Acquisition images 17, 18, and 19.

0751 - Preset for star 13 of the sequence failed, star fell short of the WFS pickoff on the stage offset.

See acquisition images 23, 24, and 25 for the sequence that failed.

0753 - Re-sent the original field preset and it acquired it just fine.

0815 - Running the transform collect again, seeing is still pretty icky...

Much fussing about, seeing is highly variable, and the IDL program was losing it well off axis.

The Missed WFS Offset Problem - Identified

In the meantime, I investigated the problem of the failed probe offsets into the WFS pickoff. These noted here and the one on the night of Nov 15 were investigated.

I found that all 3 failed presets (we had a repeat of star 13), and the near-miss (star 17 above) were, at the time of the initial acquisition, presets that had the largest Y-axis position of the guide probe before the go-to-WFS offset was executed.

The go-to-WFS offset was failing because the Y-axis stage move required for the Y-axis part of the offset was exceeding the limit of the stage travel.

Two issues:
  • The agwServer software is not sensing that the Y-axis limit sensor has been asserted and signalling a fault. A quick check with a low-level hardware query revealed that the sensor is asserting correctly, so this is a bug in the OSU software.
  • We need to modify the lower Y-axis limit of the guide patrol field to keep people from picking stars there.

The tricky bit here is that the "acquire star on guide camera field center" motion is within the valid range of motion, but the subsequent probe offset to place that star on the WFS guiding hotspot will exceed the Y-axis range of motion. This makes the logic a little non-intuitive: the first move is valid, but the subsequent move is not, but it is also not true that all probe moves are necessarily followed by a 13.x-mm offset move. The "is this move valid" test is therefore context dependent.

The easiest thing to do is limit the range of requested guide probe travel to the range of Y motion that will allow some maximum offset-to-WFS-hotspot motion. We lose a little bit of patrol field, but avoid the limit.

This will take time to implement, in the meantime we need to avoid stars near the bottom of the current "published" guide patrol field when setting up presets for transformation measurements.

Transformations redux

1027 - We have to bypass the slightly buggy IDL routine. What we will do is
  • John will send a mix of active and guide presets from IRAF by hand
  • Rick will initiate the exposure on the MODS control panel
A little loss of efficiency in a grand sense, but since it will not get hung up by the bug in the IDL WFS module, it'll come out ahead.

1113 - got 18 stars, repeated the first one as we might have issued the next preset before the exposure was done. The stack looks pretty good, a few are screwed up by the "brightness veto" imposed by the GCS in this crowded field.

What's good about Stone E is that it has lots of stars. The bad thing is that it has lots of stars...

Data are in images mods1r.20101117.0053 to 0070. Note that 0053 thru 0059 were mislabeled "Stone C", this was corrected for 60 onward.

Stars:

star Image notes
41 53 preset too soon?
45 54
106 55
55 56
68 57
24 58
1 59 ???
16 60
29 61
40 62
57 63
119 64
86 65
63 66
121 67
81 68
67 69 ??
85 70
41 71 repeat

1123 - Recentering the pointing

During acq, we had the guider camera fail.

Logged into the AzCam computer (lbt1) via VNC and found the server window listing the fault message:
ERROR Readout counter stuck
Readout aborted

We tried cycling power on the Leach controller and restarting the server program. No dice.

Then we tried rebooting the AzCam computer .. that did the trick

1139 - Trying again:

New coefficients:
  IE = 5
  CA = 23

1200 - John did a test drive of the field aberration measurement scripts in IDL, and ran into the same bug as we did in the transformation script, fixed that, and then hit a second bug that will need Doug's attention.

1215 - Seeing is still atrocious, we can't do much with this in the time remaining, so we're taking a quick recon of M67, one of our photometric reference fields

1225 - That's it, we're calling it a night.

-- RichardPogge - 17 Nov 2010
Topic revision: r10 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback